# Phantom matter: a challenging solution to the cosmological tensions

### v1

@@
Abstract. The idea of composite dark energy (DE) is quite natural since on general grounds we expect that the vacuum energy (associated to the cosmological term $\Lambda$) may appear in combination with other effective forms of DE, collectively denoted $X$. This was indeed the old proposal from 2006, (cf. Ref.[41]) called the ‘$\Lambda$XCDM model’, which was primordially designed to explain the cosmic coincidence problem. We now find that it can also have far reaching consequences on the current cosmological tensions on $H_{0}$ and the growth of large scale structure (LSS). The $\Lambda$XCDM may involve both a phantom-like component $X$ and a constant cosmological term $\Lambda$ (positive or negative) or even a running one, $\Lambda=\Lambda(H)$. In the current work, we deal with a simplified version of the model and exploit the possibility that $X$ behaves as ‘phantom matter’ (PM). The latter appears in stringy versions of the running vacuum model (RVM). Unlike phantom DE, it satisfies the strong energy condition like usual matter, hence bringing to bear positive pressure at the expense of negative energy. Bubbles of PM may appear in the manner of a transitory ‘phantom vacuum’ tunneled into the late universe before it heads towards a new de Sitter era, thereby offering a crop field for the growing of structures earlier than expected. Using SNIa, cosmic chronometers, transversal BAO, LSS data and the full CMB likelihood from Planck 2018, we find that the tensions virtually disappear in this stringy RVM scenario characterized by axionic dark matter. The value of $H_{0}$ emerging from our analysis proves compatible with SH0ES to within less than $0.25\sigma$ and the LSS growth tension is nonexistent. The statistical information criteria point to very strong evidence in favor of the PM solution.
@@ 1 Introduction
The standard (or concordance) cosmological model, aka $\Lambda$CDM, has been a rather successful paradigm for the description of the universe for more than three decades [1], especially since the late nineties [2]. Despite it constitutes the main theoretical pillar at our hands for the description of the universe’s evolution within the General Relativity (GR) context, the $\Lambda$CDM is being pestered by a number of glitches and hitches which prove to be more and more difficult to iron out. While the model is presumably right in the main (at least at a pure phenomenological level), an increasing number of worrisome inconsistencies (or “ tensions”) have been perturbing its reputation and its future prospects. For a long time the role played by a rigid cosmological term $\Lambda$ in the standard $\Lambda$CDM model has been rather successful since it has provided a fairly reasonable description of the overall cosmological data [3, 4]. But the use of the cosmological term has never been fully clarified at the formal theoretical level since it is usually associated to the so-called “cosmological constant problem” (CCP), a serious conundrum which has been amply discussed in the literature, see e.g. [5, 6, 7, 8, 9] and references therein. Inasmuch as the vacuum energy density (VED) is related to $\Lambda$ through $\rho_{\rm vac}=\Lambda/(8\pi G_{N})$ ($G_{N}$ being Newton’s gravitational coupling), the conceptual fate of $\Lambda$ is tied to our ultimate understanding of the VED on fundamental physical terms. Fortunately, recent theoretical developments on the VED in quantum field theory (QFT) are bringing new light for a potential alleviation of these theoretical difficulties. In fact, the possibility that the quantum vacuum (and hence that $\Lambda$ itself) is actually dynamical (i.e. evolving with the cosmological expansion) rather than being stuck at a rigid value, has lately been substantiated in the context of quantum field theory in curved spacetime as well as in the framework of low-energy effective string theory[8, 9, 10, 11]. For a long time, the dynamics of the VED has been popularly addressed on phenomenological grounds using ad hoc scalar fields (quintessence and the like) which supplant the role of the $\Lambda$-term through the current value of some suitable effective potential [6, 7], see e.g. [12] for a recent review. Currently, the first data release of the Dark Energy Spectroscopic Instrument (DESI) suggests tantalizing evidence that the dark energy (DE) might be dynamical using some common parameterizations of the DE [13]. More detailed analyses will be needed before getting a final confirmation of this fact, of course, but it is a fact that some anticipatory hints in the literature pointed out already this possibility a few years ago from different perspectives and using a significant amount of cosmological data. The level of evidence was substantial and ranged in between $3-4\sigma$. Some of these analysis are well-known [14, 15, 16, 17, 18] and involved the running vacuum model (RVM) [^1] For a review, see e.g. [8, 9, 19] and references therein. Other significant analysis around the same time supported also this possibility using different methods and parameterizations [20, 21].
@@
Dynamical DE could also impinge positively on the resolution of the cosmological tensions. Let us recall that these are mainly concerned with the measurement of the current Hubble parameter $H_{0}\equiv 100h$ km/s/Mpc and the growth of large scale structures (LSS). The latter is usually monitored with $S_{8}$ or $\sigma_{8}$, or even better using a parameter $\sigma_{12}$ defining the amplitude of the matter power spectrum at fixed spheres of radius $12$ Mpc rather than $8h^{-1}$ Mpc, thus avoiding artificial dependence on the value of $h$[22, 23, 24] [^2] See [25, 26] for summarized explanations about each of these tensions, and [27, 28] for comprehensive reviews.. The first kind of tension involves a serious disagreement between the CMB observations, using fiducial $\Lambda$CDM cosmology, and the local direct (distance ladder) measurements of the Hubble parameter today. It is arguably the most puzzling open question and it leads to a severe inconsistency of $\sim 5\sigma$ CL. between the mentioned observables. The second kind of tension is related with the exceeding rate of large scale structure formation predicted by the $\Lambda$CDM as compared to measurements, although the discordance here is moderate, at a confidence level of $\sim 2-3\sigma$. More recently, data from the James Webb Space Telescope (JWST) [29, 30] have revealed the existence of an unexpectedly large population of extremely massive galaxies at large redshifts $z\gtrsim 5-10$, a fact which is also strongly at odds with the prospects of the concordance model.
@@
There is a wide panoply of strategies in the literature trying to mitigate some of the above tensions, although at present we may be still far away from a satisfactory resolution of the situation. We mention only a few. It has been argued that within the class of models where the DE is dealt with as a cosmic fluid with equation of state (EoS) $w(z)$, solving the $H_{0}$ tension demands the phantom condition $w(z)<-1$ at some $z$, whereas solving both the $H_{0}$ and $\sigma_{8}$ tensions requires $w(z)$ to cross the phantom divide and/or other sorts of exotic transitions, see e.g. [31, 32, 33, 34, 35, 36, 37]. The possibility of a sign flip of the $\Lambda$ term has been entertained in recent times, e.g. in [38]. Let us also mention the model analyzed in [39] (based on the framework of [40]), in which one considers a sudden transition from anti-de Sitter (AdS), hence $\Lambda<0$, into de Sitter (dS) regime ($\Lambda>0$) occurring near our time. Despite it being essentially ad hoc, the model yields a rather good fit to the data when transversal BAO is employed in the fitting analysis, see also [37]. A more general framework already existed since long ago, namely the $\Lambda$XCDM model [41, 42, 43], where the running vacuum and an extra $X$ component can exchange energy. In this context, one can have $\Lambda<0$ or $\Lambda>0$.
@@
In this work, we further exploit the virtues of the $\Lambda$XCDM class, which is an enhanced version of the RVM approach [8]. The linchpin of the RVM framework is the dynamical nature of the VED framed in the fundamental context of QFT in curved spacetime or in low-energy string theory. Appropriate renormalization in this context gets rid of the quartic mass terms $\sim m^{4}$ responsible for the fine tuning troubles, and as a result the VED becomes a smooth function of even powers of the Hubble rate [44, 45, 46, 47]. This leads to a fairly good description of the cosmological data and helps easing the tensions [48, 49]. Worth noticing is also the phenomenological performance of the stringy version of the RVM[50]. All that said, we cannot exclude that the DE may be a composite fluid, a possibility which could further help in the task of alleviating these tensions. For this reason a good candidate is the old existing ‘$\Lambda$XCDM model’ [41, 42, 43], a RVM-born model which was initially motivated as a possible cure for the cosmic coincidence problem. In it, we have two DE components: one is the running VED, $\rho_{\rm vac}=\rho_{\rm vac}(H)$, and the other is the mentioned $X$. As a matter of fact, the entity $X$ need not be a fundamental field, as emphasized in [41]: it may just be due to particular terms in the effective action which mimic a dynamical DE component. In this sense, $X$ can have phantom-like behavior without causing uproar at the theoretical level. Not only so, $X$ may even display “phantom matter” (PM) behavior[41], namely an intriguing form of DE which, in contrast to the usual phantom DE, is characterized by a positive pressure ($p_{X}>0$) at the expense of a negative energy density ($\rho_{X}<0$). It is remarkable that there exists specific theoretical scenarios in the current literature which support the $\Lambda$XCDM structure with a PM-like component $X$. An example appears in the stringy RVM context [10, 11]. No less remarkable is the fact that the phenomenological performance attained by such a peculiar composite DE scenario surpasses by far that of the standard $\Lambda$CDM, as we shall shortly report. However, for the sake of simplicity, in this work we shall address a reduced version of the $\Lambda$XCDM, which we call the $w$XCDM model, in which the $X$ component of the former is kept intact, but on the other hand we mimic the running $\Lambda$ with another dynamical component which we call $Y$. While the analysis within the full $\Lambda$XCDM will be presented elsewhere, we shall nonetheless demonstrate here that even this simplified $w$XCDM version of the $\Lambda$XCDM proves extremely efficient in dealing with the cosmological tensions, provided that the $X$ component behaves as phantom matter. The information criteria applied to the results of our fitting analyses indeed bolster such an option as the most competitive one (by far). Therefore, we have a good theoretical framework to accommodate the phantom matter scenario, which proves phenomenologically optimal.
@@ 2 Composite dark energy
We next mention three related types of composite models of the DE: i) $\Lambda$XCDM, ii) $w$XCDM and iii) $\Lambda_{s}{\rm CDM}$. As noted, model i) exists in different versions since long ago [41, 42, 43]; model ii) will be analyzed here for the first time, it constitutes a simplified version of model i) and embodies the PM feature. Finally, model iii) was recently analyzed in [39], see also references therein.
$\bullet$ i) $\Lambda$XCDM model. Its definition and comprehensive discussion, including the background cosmological solution, are provided in utmost detail in [41]. In addition, in [43] the corresponding cosmic perturbations equations are fully accounted for. Here we just present a very short qualitative description. Within the $\Lambda$XCDM, the cosmic fluid contains the usual matter energy density $\rho_{m}$ and a composite DE sector made out of two components, one is the running vacuum energy density $\rho_{\rm vac}$ and the other is called $X$, with energy density $\rho_{X}$. The VED here is treated within the QFT framework of the RVM [44, 45, 46, 47], in which $\rho_{\rm vac}=\rho_{\rm vac}(H)$ evolves with the expansion rate. In the current universe, such an evolution reads [8]:
@@
with $\rho^{0}_{\rm vac}$ the current VED value, $m_{\rm Pl}$ the Planck mass and $|\nu_{\rm eff}|\ll 1$ a small parameter which is formally computable in QFT, see the aforementioned papers. For $\nu_{\rm eff}>0$ the VED decreases with the expansion and hence the RVM mimics quintessence, whereas for $\nu_{\rm eff}<0$ the VED increases with the expansion and the RVM behaves effectively as phantom DE. The measured cosmological term is $\Lambda=8\pi G\rho_{\rm vac}$, and hence $\Lambda$ also runs with $H$ since $\rho_{\rm vac}=\rho_{\rm vac}(H)$. Such a running feature of the VED and $\Lambda$ occurs in the $\Lambda$XCDM too, and in exactly the same form (1), but here we have also the dynamics of $X$ and hence the cosmological solution of the model in terms of the redshift variable is considerably more complicated [43]. Furthermore, as recently demonstrated in the QFT context, $\rho_{\rm vac}$ has an EoS which departs from $-1$[46], so the “modern version of the $\Lambda$XCDM model” has actually two nontrivial EoS parameters, one for the running VED and the other for $X$, which we call $\omega_{X}$. These are basic ingredients in our analysis, and for this reason the $\Lambda$XCDM fleshes out the theoretical basis inspiring the current work. However, the actual scenario which we will analyze here is model $w$XCDM (see below). It is simpler than the $\Lambda$XCDM but it has the same genetic ingredients. We will use it to emulate the basic properties of the latter. The analysis of the cosmic tensions within the full $\Lambda$XCDM framework is more demanding and will be presented in a separate work[51].
@@
$\bullet$ ii) $w$XCDM model. In order to illustrate the possibilities of our composite DE scenario in connection with the cosmological tensions, herein we will use the $w$XCDM model, a reduced (skeleton) version of the $\Lambda$XCDM. The two models share the $X$ component, but $w$XCDM mimics the running vacuum feature of $\Lambda$XCDM through another dynamical component $Y$ (replacing $\Lambda$) and whose EoS $w_{Y}$ can be different from $-1$. This means that $Y$ does not act as a rigid cosmological constant, which is fair enough since, as mentioned before, in the RVM context the EoS of the VED (hence of $\Lambda$) actually departs from $-1$ owing to quantum effects [46]. Thus, in the $w$XCDM , we have two DE components with different EoS behavior, to wit: the $X$ component, which behaves phantom-like ($\omega_{X}\lesssim-1$) and the $Y$ component behaving quintessence-like ($\omega_{Y}\gtrsim-1$). An important point is that $X$ and $Y$ do not act simultaneously: $X$ acts first in the cosmic expansion, whilst $Y$ acts subsequently. To be precise, $X$ enters only above a transition redshift $z>z_{t}$ (fitted from the data) and $Y$ enters below that redshift until the current time. No less crucial is the fact that the phantom-like component $X$ actually behaves as PM, therefore with negative energy density ($\Omega_{X}=\rho_{X}/\rho_{c}<0$) and positive pressure ($p_{X}>0$), whereas for the $Y$ component we have $\Omega_{Y}>0$, $p_{Y}<0$. Notice that the characteristic free parameters of model $w$XCDM are just $(z_{t},w_{X},w_{Y})$. Indeed, the density parameters for the DE components are not free since e.g. the value of $\Omega_{Y}^{0}\equiv\Omega_{Y}(z=0)$ depends on the fitting values of $H_{0},\omega_{b},\omega_{\rm dm}$ (cf. Table 1). In addition, the respective values of $\Omega_{X}$ and of $\Omega_{Y}$ immediately above and below $z_{t}$ are assumed to be equal in absolute value. Thus, since we find $\Omega_{X}<0$ and $\Omega_{Y}>0$, this means that $|\Omega_{X}(z)|=\Omega_{Y}(z)$ at $z=z_{t}$. This kind of assumption aims at reducing the number of parameters and is entirely similar to the one made in the $\Lambda_{s}{\rm CDM}$ model, see point iii) below.
@@
Last, but not least, we note that the existence of these two different phases of the DE separated by the redshift $z_{t}$ is not just an ad hoc assumption since it is found in theoretical contexts such as the stringy RVM approach [11], which point to the existence of transitory domains or bubbles of PM whenever the universe is approaching a de Sitter epoch (see next sections for more elaboration).
@@
$\bullet$ iii) $\Lambda_{s}{\rm CDM}$ model. This is the model recently analyzed with considerable phenomenological success in [39]. Strictly speaking, it is not a composite model since only $\Lambda$ is involved, although it enters with two signs and in this sense it is a composition of two phases of $\Lambda$ separated also by a transition redshift $z_{t}$, which is the characteristic parameter of this model. It is assumed that at that point there is a sudden AdS-dS transition from $-\Lambda<0$ in the upper range to $+\Lambda>0$ in the lower range, i.e. an abrupt change of sign of $\Lambda$ at $z=z_{t}$, but keeping the same absolute value. We refer the reader to the quoted reference for more details.
@@
We shall simultaneously provide here the fitting results for models ii) and iii) under the very same data sources. Together with the standard $\Lambda$CDM model, the simultaneous analysis of $\Lambda_{s}{\rm CDM}$ will be useful as a benchmark for rating the performance of the phantom matter approach $w$XCDM to the cosmological tensions. At the same time we will provide the results for the standard $\Lambda$CDM.

### v3

@@
Abstract
The idea of composite dark energy (DE) is quite natural since on general grounds we expect that the vacuum energy density (associated with the cosmological term $\Lambda$) may appear in combination with other effective forms of DE, denoted $X$. Here we deal with model $w$XCDM, a simplified version of the old $\Lambda$XCDM model (Grande et al. (2006)), and exploit the possibility that $X$ behaves as ‘phantom matter’ (PM), which appears in stringy versions of the running vacuum model (RVM). Unlike phantom DE, the PM fluid satisfies the strong energy condition like usual matter, hence bringing to bear positive pressure at the expense of negative energy. Bubbles of PM may appear in the manner of a transitory ‘phantom vacuum’ tunneled into the late universe before it heads towards a new de Sitter era, thereby offering a crop field for the growing of structures earlier than expected. Using SNIa, cosmic chronometers, transversal BAO (BAO 2D), LSS data and the full CMB likelihood from Planck 2018, we find that the $H_{0}$ and growth tensions virtually disappear, provided that BAO 2D are the only source of BAO data used in the fit. In contrast, our preliminary analysis using exclusively anisotropic BAO (BAO 3D) indicates that the ability to ease the $H_{0}$ tension is significantly reduced as compared to the scenario with BAO 2D, despite the fact that the overall fit to the cosmological data is still better than in the $\Lambda$CDM. Finally, our approach with BAO 2D favors quintessence-like behavior of the DE below $z\simeq 1.5$ at $\gtrsim 3\sigma$ CL, which is compatible with the recent DESI measurements.
stars: Cosmology — Cosmological models — Cosmological parameters – Cosmological evolution – Dark energy – Large-scale structure of the universe
@@ 1 Introduction
The standard (or concordance) cosmological model, aka $\Lambda$CDM, has been a rather successful paradigm for the description of the universe for more than three decades (Peebles, 1993), especially since the late nineties (Turner, 2022). Despite it constitutes the main theoretical pillar at our hands for the description of the universe’s evolution within the General Relativity (GR) context, the $\Lambda$CDM is being pestered by a number of glitches and hitches which prove to be more and more difficult to iron out. While the model is presumably right in the main (at least at a pure phenomenological level), an increasing number of worrisome inconsistencies (or “ tensions”) have been perturbing its reputation and its future prospects. For a long time the role played by a rigid cosmological constant (CC) $\Lambda$ in the standard $\Lambda$CDM model has been rather successful since it has provided a fairly reasonable description of the overall cosmological data (Ade et al., 2016; Aghanim et al., 2020). But the use of the cosmological term has never been fully clarified at the formal theoretical level since it is usually associated to the so-called “cosmological constant problem”, a serious conundrum which has been amply discussed in the literature, see e.g. (Weinberg, 1989; Peebles & Ratra, 2003; Padmanabhan, 2003; Solà Peracaula, 2022; Solà, 2013) and references therein. Inasmuch as the vacuum energy density (VED) is related to $\Lambda$ through $\rho_{\rm vac}=\Lambda/(8\pi G_{N})$ ($G_{N}$ being Newton’s gravitational coupling), the conceptual fate of $\Lambda$ is tied to our ultimate understanding of the VED on fundamental physical terms. Fortunately, recent theoretical developments on the VED in quantum field theory (QFT) are bringing new light for a potential alleviation of these theoretical difficulties. In fact, the possibility that the quantum vacuum (and hence that $\Lambda$ itself) is actually dynamical (i.e. evolving with the cosmological expansion) rather than being stuck at a rigid value, has lately been substantiated in the context of QFT in curved spacetime (Solà Peracaula, 2022; Solà, 2013) as well as in the framework of low-energy effective string theory (Mavromatos & Solà Peracaula, 2021a, b) and, very recently, in lattice quantum gravity (Dai et al., 2024). For a long time, the dynamics of the VED has been popularly addressed on phenomenological grounds using ad hoc scalar fields (quintessence and the like) which supplant the role of the $\Lambda$-term through the current value of some suitable effective potential (Peebles & Ratra, 2003; Padmanabhan, 2003), see e.g. (Avsajanishvili et al., 2024) for a recent review. Currently, the first data release of the Dark Energy Spectroscopic Instrument (DESI) suggests tantalizing evidence that the dark energy (DE) might be dynamical using some common parameterizations of the DE (Adame et al., 2024). More detailed analyses will be needed before getting a final confirmation, of course, but it is a fact that some anticipatory (and fairly robust) hints of dynamical DE in the literature already pointed out this possibility a few years ago from different perspectives and using a significant amount of cosmological data. The level of evidence put forward by these anticipatory studies was substantial and ranged $3-4\sigma$. Some of these analysis are well-known (Solà et al., 2015, 2017; Solà Peracaula et al., 2017, 2018a, 2018b; Gómez-Valent & Solà Peracaula, 2018) and involved the running vacuum model (RVM) [^1] For a review, see e.g. (Solà Peracaula, 2022; Solà, 2013; Solà & Gómez-Valent, 2015) and references therein.. Subsequent analysis around the same time supported also this possibility using different methods and parameterizations (Zhao et al., 2017; Solà Peracaula et al., 2019).
@@
Dynamical DE could also impinge positively on the resolution of the cosmological tensions. Let us recall that these are mainly concerned with the measurement of the current Hubble parameter $H_{0}\equiv 100h$ km/s/Mpc and the growth of large scale structures (LSS). The latter is usually monitored with $S_{8}$ or $\sigma_{8}$, or even better using a parameter $\sigma_{12}$ defining the amplitude of the matter power spectrum at fixed spheres of radius $12$ Mpc rather than $8h^{-1}$ Mpc, thus avoiding artificial dependence on the value of $h$(Sánchez, 2020; Semenaite et al., 2022, 2023) [^2] See (Di Valentino et al., 2021a, b) for summarized explanations about each of these tensions, and (Perivolaropoulos & Skara, 2022a; Abdalla et al., 2022) for comprehensive reviews.. The first kind of tension involves a serious disagreement between the CMB observations, using fiducial $\Lambda$CDM cosmology, and the local direct (distance ladder) measurements of the Hubble parameter today. It is arguably the most puzzling open question and it leads to a severe inconsistency of $\sim 5\sigma$ c.l. between the mentioned observables. The second kind of tension is related with the exceeding rate of large scale structure formation in the late universe predicted by the $\Lambda$CDM as compared to measurements, although the discordance here is moderate, at a confidence level of $\sim 2-3\sigma$. More recently, data from the James Webb Space Telescope (JWST) (Gardner et al., 2006; Labbé et al., 2023) have revealed the existence of an unexpectedly large population of extremely massive galaxies at large redshifts $z\gtrsim 5-10$, a fact which is also strongly at odds with the prospects of the concordance model.
@@
There is a wide panoply of strategies in the literature trying to mitigate some of the above tensions, although at present we may be still far away from a satisfactory resolution of the situation. We mention only a few. It has been argued that within the class of models where the DE is dealt with as a cosmic fluid with equation of state (EoS) $w(z)$, solving the $H_{0}$ tension demands the phantom condition $w(z)<-1$ at some $z$, whereas solving both the $H_{0}$ and $\sigma_{8}$ tensions requires $w(z)$ to cross the phantom divide and/or other sorts of exotic transitions, see e.g. (Heisenberg et al., 2022; Marra & Perivolaropoulos, 2021; Alestas et al., 2021; Perivolaropoulos & Skara, 2021; Alestas et al., 2022; Perivolaropoulos & Skara, 2022b; Gómez-Valent et al., 2024). The possibility of a sign flip of the $\Lambda$ term has been entertained in recent times, e.g. in (Calderón et al., 2021). Let us also mention the model analyzed in (Akarsu et al., 2021, 2023) (based on the framework of Akarsu et al. (2020)), in which one considers a sudden transition from anti-de Sitter (AdS), hence $\Lambda<0$, into de Sitter (dS) regime ($\Lambda>0$) occurring near our time. Despite it being essentially ad hoc, the model yields a rather good fit to the data when transversal/angular BAO (BAO 2D for short) is employed in the fitting analysis, see also (Gómez-Valent et al., 2024). Actually, a more general framework already existed since long ago, namely the $\Lambda$XCDM model (Grande et al., 2006, 2007, 2009), where the running vacuum and an extra $X$ component can exchange energy. In such a context, one can have $\Lambda<0$ or $\Lambda>0$.
@@
In this work, we further exploit the virtues of the $\Lambda$XCDM, which is an enhanced family of RVM’s (Solà Peracaula, 2022). The linchpin of the RVM framework is the dynamical nature of the VED framed in the fundamental context of QFT in curved spacetime or in low-energy string theory. Appropriate renormalization in this context gets rid of the quartic mass terms $\sim m^{4}$ responsible for the fine tuning troubles, and as a result the VED becomes a smooth function of even powers of the Hubble rate (Moreno-Pulido & Solà, 2020; Moreno-Pulido & Solà Peracaula, 2022, 2022; Moreno-Pulido et al., 2023). This leads to a fairly good description of the cosmological data and helps easing the tensions (Solà Peracaula et al., 2021, 2023). Worth noticing is also the phenomenological performance of the stringy version of the RVM (Gómez-Valent et al., 2023). This said, we cannot exclude that the DE may be a composite fluid, a possibility which could further help in the task of alleviating these tensions. For this reason a good candidate is the aforementioned ‘$\Lambda$XCDM model’ (Grande et al., 2006, 2007, 2009), a RVM-born model which was initially motivated as a possible cure for the cosmic coincidence problem. In it, we have two DE components: one is the running VED, $\rho_{\rm vac}=\rho_{\rm vac}(H)$, and the other is the mentioned $X$. As a matter of fact, the entity $X$ need not be a fundamental field, as emphasized in (Grande et al., 2006): it may just be due to particular terms in the effective action which mimic a dynamical DE component. In this sense, $X$ can have phantom-like behavior without causing uproar at the theoretical level. Not only so, $X$ may even display “phantom matter” (PM) behavior (Grande et al., 2006), namely an intriguing form of DE which, in stark contrast to the usual phantom DE, is characterized by a positive pressure ($p_{X}>0$) at the expense of a negative energy density ($\rho_{X}<0$). It is remarkable that there exists specific theoretical scenarios in the current literature which support the $\Lambda$XCDM structure with a PM-like component $X$. An example appears in the stringy RVM context (Mavromatos & Solà Peracaula, 2021a, b). No less remarkable is the fact that the phenomenological performance attained by such a peculiar composite DE scenario may surpass by far that of the standard $\Lambda$CDM under appropriate conditions, as we shall shortly clarify. However, for the sake of simplicity, in this paper we shall address a reduced version of the $\Lambda$XCDM, which we call the $w$XCDM model (not to be confused with the conventional $w$CDM parameterization), in which the additional $X$ component of the $\Lambda$XCDM is kept intact, but on the other hand we mimic the running $\Lambda$ with another dynamical component which we call $Y$. While the analysis within the full $\Lambda$XCDM will be presented elsewhere, we shall nonetheless demonstrate here that even the simplified $w$XCDM model can be very efficient in dealing with the cosmological tensions. However, for this to be so the following two conditions must be met in our analysis: i) the $X$ component must have the ability to behave as phantom matter; and ii) BAO 2D must be the only source of BAO data used in our global fit. The precise reasons for considering only this type of BAO data in the current work will be explained below, together with a preliminary discussion of the results obtained with the BAO 3D variant. The upshot is that by accepting these reasons for using BAO 2D only, and also taking advantage of the mentioned (stringy) theoretical framework to accommodate phantom matter, the possibility to quell the cosmological tensions on fundamental grounds is viable and can be extremely effective.
@@ 2 Composite dark energy
We next mention three related types of composite models of the DE: i) $\Lambda$XCDM, ii) $w$XCDM and iii) $\Lambda_{s}{\rm CDM}$. As noted, model i) exists in different versions since long ago (Grande et al., 2006, 2007, 2009); model ii) will be analyzed here for the first time, it constitutes a simplified version of model i) and embodies the PM feature. Finally, model iii) was recently analyzed in (Akarsu et al., 2023).
$\bullet$ i) $\Lambda$XCDM model. Its definition and comprehensive discussion, including the background cosmological solution, are provided in utmost detail in (Grande et al., 2006). In addition, in (Grande et al., 2009) the corresponding cosmic perturbations equations are fully accounted for. Here we just present a very short qualitative description. Within the $\Lambda$XCDM, the cosmic fluid contains the usual matter energy density $\rho_{m}$ and a composite DE sector made out of two components, one is the running vacuum energy density $\rho_{\rm vac}$ and the other is called $X$, with energy density $\rho_{X}$. The VED here is treated within the QFT framework of the RVM (Moreno-Pulido & Solà, 2020; Moreno-Pulido & Solà Peracaula, 2022, 2022; Moreno-Pulido et al., 2023), in which $\rho_{\rm vac}=\rho_{\rm vac}(H)$ evolves with the expansion rate. In the current universe, such an evolution reads (Solà Peracaula, 2022):
@@
with $\rho^{0}_{\rm vac}$ the current VED value, $m_{\rm Pl}$ the Planck mass and $|\nu_{\rm eff}|\ll 1$ a small parameter which is formally computable in QFT, see the aforementioned papers. For $\nu_{\rm eff}>0$ the VED decreases with the expansion and hence the RVM mimics quintessence, whereas for $\nu_{\rm eff}<0$ the VED increases with the expansion and the RVM behaves effectively as phantom DE. The measured cosmological term is $\Lambda=8\pi G\rho_{\rm vac}$, and hence $\Lambda$ also runs with $H$ since $\rho_{\rm vac}=\rho_{\rm vac}(H)$. Such a running feature of the VED and $\Lambda$ occurs in the $\Lambda$XCDM too, and in exactly the same form (1), but here we have also the dynamics of $X$ and hence the cosmological solution of the model in terms of the redshift variable is considerably more complicated (Grande et al., 2006). Besides, as recently demonstrated in the QFT context, $\rho_{\rm vac}$ has an EoS which departs from $-1$ (Moreno-Pulido & Solà Peracaula, 2022), so the “modern version of the $\Lambda$XCDM model” has actually two nontrivial EoS parameters, one for the running VED and the other for $X$, which we call $\omega_{X}$. These are basic ingredients in our analysis, and for this reason the $\Lambda$XCDM fleshes out the theoretical basis inspiring the current work. However, the actual scenario which we will analyze here is model $w$XCDM (see below). It is simpler than the $\Lambda$XCDM but it has the same genetic ingredients. We will use it to emulate the basic properties of the latter. The analysis of the cosmic tensions within the full $\Lambda$XCDM framework is more demanding and will be presented in a separate work (Gómez-Valent & Solà Peracaula, 2024).
@@
$\bullet$ ii) $w$XCDM model. In order to illustrate the possibilities of our composite DE scenario in connection with the cosmological tensions, herein we will use the $w$XCDM model, a reduced (skeleton) version of the $\Lambda$XCDM. The two models share the $X$ component, but $w$XCDM mimics the running vacuum feature of $\Lambda$XCDM through another dynamical component $Y$ (replacing $\Lambda$) and whose EoS, $w_{Y}$, can be different from $-1$. This means that $Y$ does not act as a rigid CC, which is fair enough since, as mentioned before, in the RVM context the VED is dynamical and its EoS (hence that of $\Lambda$) actually departs from $-1$ owing to quantum effects (Moreno-Pulido & Solà Peracaula, 2022). Thus, in the $w$XCDM , we have two DE components with potentially different EoS behaviors. Furthermore, as we will see in section 5, in the best-fit model the $X$ component behaves phantom-like ($\omega_{X}\lesssim-1$) and the $Y$ component behaves quintessence-like ($\omega_{Y}\gtrsim-1$). An important point is that $X$ and $Y$ do not act simultaneously: $X$ acts first in the cosmic expansion, and $Y$ acts subsequently. To be precise, $X$ enters only above a transition redshift $z>z_{t}$ (fitted from the data) whilst $Y$ enters below that redshift until the current time. No less crucial is the fact that the phantom-like component $X$ actually behaves as PM, therefore with negative energy density ($\Omega_{X}=\rho_{X}/\rho_{c}<0$) and positive pressure ($p_{X}>0$), whereas for the $Y$ component we have $\Omega_{Y}>0$, $p_{Y}<0$. Notice that the characteristic free parameters of model $w$XCDM are just $(z_{t},w_{X},w_{Y})$. Indeed, the density parameters for the DE components are not free since e.g. the value of $\Omega_{Y}^{0}\equiv\Omega_{Y}(z=0)$ depends on the fitting values of $H_{0},\omega_{b},\omega_{\rm dm}$ (cf. Table 1). In addition, the respective values of $\Omega_{X}$ and of $\Omega_{Y}$ immediately above and below $z_{t}$ are assumed to be equal in absolute value. This means that $|\Omega_{X}(z)|=\Omega_{Y}(z)$ at $z=z_{t}$. This kind of assumption aims at reducing the number of parameters and is entirely similar to the one made in the $\Lambda_{s}{\rm CDM}$ model, see point iii) below.
@@
Last, but not least, we note that the existence of these two different phases of the DE separated by the redshift $z_{t}$ is not just an ad hoc assumption since it is found in theoretical contexts such as the stringy RVM approach (Mavromatos & Solà Peracaula, 2021b), which points to the existence of transitory domains or bubbles of PM whenever the universe is approaching a de Sitter epoch (see next sections for more elaboration).
@@
$\bullet$ iii) $\Lambda_{s}{\rm CDM}$ model. This is the model recently analyzed with considerable phenomenological success in (Akarsu et al., 2023). Strictly speaking, it is not a composite model since only $\Lambda$ is involved, although it enters with two signs and in this sense it is a composition of two phases of $\Lambda$ separated also by a transition redshift $z_{t}$, which is the characteristic parameter of this model. It is assumed that at that point there is a sudden AdS-dS transition from $-\Lambda<0$ in the upper redshift range to $+\Lambda>0$ in the lower range, i.e. an abrupt change of sign of $\Lambda$ at $z=z_{t}$, but keeping the same absolute value. We refer the reader to the quoted reference for more details.
@@
For convenience, we shall simultaneously provide the fitting results for models ii) and iii) under the very same data sources. Together with the standard $\Lambda$CDM model, the simultaneous analysis of $\Lambda_{s}{\rm CDM}$ will be useful as a benchmark for rating the performance of the phantom matter approach $w$XCDM to the cosmological tensions.